OpenStudy (anonymous):

Calculating Molarity: A student needs to make 0.755 L of a sodium bromide solution for an experiment. The concentration of the required solution is 0.450 M. How many grams of sodium bromide will she need to use?

imqwerty (imqwerty):

okay so do you know the formula to find molarity? :)

imqwerty (imqwerty):

@chemhelppleasees

OpenStudy (anonymous):

M = mol / liter?

imqwerty (imqwerty):

yeah :) \(\large M = \Large \frac{n}{L}\) here \(M=molarity\) \(n=number~of~moles\) \(L=volume~in~liters\) okay so now tell how do we calculate number of moles

OpenStudy (anonymous):

I'm not sure but is this how you would set it up \[0.450 = x/0.755\]?

imqwerty (imqwerty):

yeah thats correct can you find \(x\) from here

OpenStudy (anonymous):

x = .33975

imqwerty (imqwerty):

yes correct now we have the number of moles=0.33975 now we gotta find mass of \(NaBr\) used

OpenStudy (anonymous):

If the molar mass of NaBr is 102.894 would you multiply that by .33975 to get 34.958?

imqwerty (imqwerty):

try to use this formula- \(\large moles= \Large \frac{mass}{Molar~Mass}\) you know the number of moles Molar mass of \(NaBr\) is 102.89 just put them to get the mass of \(NaBr\)

imqwerty (imqwerty):

yea you are correct you gotta multiply them up and your answer is also correct :)

OpenStudy (anonymous):

Okay, thank you for your help!

imqwerty (imqwerty):

np :)